OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is parallel to the line y = 3x +2 and passes through the point (5, 2). A. y = 3x – 5 B. y = 3x – 13 C. y = 3x – 1 D. y = 3x + 2

OpenStudy (anonymous):

If its parallel to y=3x+2 that means it has the same slope, can you pick out what the slope is from that equation?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

No, the slope is 3. The equation is in the form y=mx+b where m is slope and b is the y intercept. Parallel lines have the same slope. so since we have the slope now and a point, we can do slope point form y-2=3(x-5) y-2=3x-15 y=3x-13